• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 08:38:48 on January26,2002

    Entry number 80858

    keyword=shift summary

    0:00-0:16 4 pass beam tuning in Hall A
    0:30 access for Q1 reset
    1:20 the Q1 trip problem has been solved
    2:10 optics run for kinematics 4H
    2:55 IOC reboot when moving the target
    data for 4H
    4:10 lost beam
    5:20 beam, 18uA continue 4H
    6:25 IOC reboot when moving to the LH2 R target
    7:00 lost TS2 connection, reset CODA
    7:50 lost TS2, ROC4 connection, again CODA reset